# GuRE:Generative Query REwriter for Legal Passage Retrieval

### v1

@@ 2 Method: GuRE
Table 1: Evaluation results for various retrieval methods with different numbers of target passages ($N$k). The best performance for each retriever, across all metrics, is highlighted in bold. $\dagger$ denotes a statistically significant improvement (paired $t$-test, $p<0.01$) over the best-performing method excluding those marked in bold.
@@ 4.1 Main Results
In contrast, other baseline methods yield suboptimal performance gains, falling short of the improvements by GuRE. Q2D achieves the lowest performance gain, suggesting that the few-shot prompting strategy struggles to address the underlying challenges in tasks requiring domain-specific knowledge. Furthermore, retriever fine-tuning does not provide retrievers with the same level of performance as GuRE. This indicates that mitigating vocabulary mismatch through query rewriting with GuRE is significantly more effective than training the retrievers.
@@
Table 2: Quantitative evaluation of pseudo-passages (Q2D, Q2D-CoT) and "rewritten query" (GuRE) between target passages on the 10K test set. The best values for each column are highlited in bold.
@@ 4.2 Rewritten Query Evaluation
Additionally, we find that the "rewritten query" generated by GuRE contains semantically similar legal context to the target passage (Table 3). For example, GuRE successfully generates phrases like "action for trademark infringement". In contrast, pseudo-passages from Q2D are mostly irrelevant, and while Q2D-CoT generates some relevant context like "trademark infringement", it also produces irrelevant context such as "defendant’s intent in adopting its mark". These results show that training an LLM with domain-specific knowledge is far more effective than relying on few-shot prompting for mitigating vocabulary mismatch. More case-studies are in the appendix I.
@@ 4.3 Which Model Should We Train?
Citations in U.S. federal precedents follow a long-tailed distribution, with the top 1% of passages accounting for 18% of all citations, while 64% receive only one citation Mahari et al. (2024). To investigate the impact of this imbalance, we analyze performance changes by varying the frequency thresholds of test samples. We sort test candidates (10%) by their frequency in the training set (90%) and select from the top Xmost frequent passages (X = 10, 30, 50, 70, 90) from test candidates. As X increases, the test set includes more unique passages. We sample 10,000 examples per threshold.
@@ Appendix A Details of Baselines
[…]
@@
[…]
@@ C.1 Prompts
Table 5 compares in-context example selection methods. While Q2D-TOP3 uses pseudo-relevant examples, its advantage is limited to R@@@@10, suggesting that example selection methods do not significantly impact performance. So, we use fixed random examples following Wang et al. (2023).
@@ Appendix F Evaluation Metrics
We evaluate the performance of our retrievers using Recall@@@@1, Recall@@@@10, nDCG@@@@10. Recall@@@@1 measures the proportion of queries for which the correct passage is ranked first in the retrieved list. Recall@@@@10 extends this by measuring the proportion of queries for which the correct passage appears in the top 10 retrieved passages. It reflects the model’s ability to identify relevant passages within a broader set of candidates. nDCG@@@@10 (Normalized Discounted Cumulative Gain at 10) considers the position of relevant passages, giving higher weight to passages ranked closer to the top.
@@ Appendix I Case Studies
Other baseline methods struggle to retrieve the target passage due to vocabulary mismatches between the query and the target passage (Table 8, 11) or because the expanded query includes irrelevant information which may incur hallucination problems mentioned in Introduction (Table 9, 10). However, GuRE generates a query identical to the target passage (Table 12).

### v2

@@ 2 Method: GuRE
Table 1: Evaluation results for various retrieval methods with different numbers of target passages (NNk). The best performance for each retriever, across all metrics, is highlighted in bold. $\dagger$ denotes a statistically significant improvement (paired $t$-test, $p<0.01$) over the best-performing method excluding those marked in bold.
@@ 4 Results
In contrast, other baseline methods yield suboptimal performance gains, falling short of the improvements by GuRE. Q2D achieves the lowest performance gain, suggesting that the few-shot prompting strategy struggles to address the underlying challenges in tasks requiring domain-specific knowledge. Furthermore, retriever fine-tuning does not provide retrievers with the same level of performance as GuRE. This indicates that mitigating vocabulary mismatch is significantly more effective than training the retrievers.
@@
Table 2: Quantitative evaluation of pseudo-passages (Q2D, Q2D-CoT) and "rewritten query" (GuRE) between target passages on the 10K test set.
@@ 5.1 Rewritten Query Evaluation
Additionally, we find that the "rewritten query" generated by GuRE contains semantically similar legal context to the target passage (Table 3). For example, GuRE successfully generates phrases like "action for trademark infringement". In contrast, pseudo-passages from Q2D are mostly irrelevant, and while Q2D-CoT generates some relevant context like "trademark infringement", it also produces irrelevant context such as "defendant’s intent in adopting its mark". These results show that domain-specific training outperforms few-shot prompting in mitigating vocabulary mismatch. More case-studies are in the appendix I.
10K Cases
R@1 R@10 nDCG@10
ModernBert 7.11 22.47 13.94
GuRE (10K) + ModerBert 16.42 39.02 26.58
GuRE (100 K) + ModernBert 20.62 45.98 32.31
20K Cases
ModernBert 6.04 19.16 11.9
GuRE (10K) + ModerBert 12.06 31.29 20.81
GuRE (100 K) + ModernBert 15.35 37.09 24.46
50K Cases
ModernBert 4.94 15.24 9.58
GuRE (10K) + ModerBert 8.67 23.94 15.53
GuRE (100 K) + ModernBert 10.3 26.66 17.71
Table 4: Retrieval results of GuRE trained under data-scarce settings. GuRE with only 10K training examples outperforms retriever fine-tuning approaches that require millions of examples across all retrieval pools.
5.2 Generalizability under Data Constraints
Although GuRE is designed as a plug-and-play, retriever-agnostic approach, it still requires training. To assess its applicability in data-scarce environments, such as legal systems where case law is only partially available, we conducted experiments with varying training sizes. Results show that GuRE trained on only 10K cases already outperforms retriever fine-tuning across all retrieval pool settings. When trained on 100K cases—a scale more realistic for practical deployment—performance further improves. These findings demonstrate that GuRE remains robust under limited-resource conditions and holds strong potential for practical use across diverse legal systems.
@@ 5.3 Which Model Should We Train?
Citations in U.S. federal precedents follow a long-tailed distribution, with the top 1% of passages accounting for 18% of all citations, while 64% receive only one citation Mahari et al. (2024). To investigate the impact of this imbalance, we analyze performance changes by varying the frequency thresholds of test samples. We sort test candidates (10%) by their frequency in the training set (90%) and select from the top X% most frequent passages (X = 10, 30, 50, 70, 90) from test candidates. As X increases, the test set includes more unique passages. We sample 10,000 examples per threshold.
@@ 6 Conclusion
Acknowledgments
This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2025-RS-2020-II201789) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation, Contribution Rate: 45%). This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2025 (Project Name: Development of an AI-Based Korean Diagnostic System for Efficient Korean Speaking Learning by Foreigners, Project Number: RS-2025-02413038, Contribution Rate: 45%). This work was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH), Contribution Rate: 10%).
@@ Appendix A Details of Baselines
$\text{LLM(Prompt}_{q}\text{)}$ represent generated pseudo passage from few-shot Q2D prompt. Q2D uses $q^{+}$ to retrieve the most relevant passage.
@@
$\text{LLM(CoTPrompt}_{q}\text{)}$ represent generated pseudo passage from few-shot Q2D-CoT prompt. Q2D-CoT uses $q^{+}$ to retrieve the most relevant passage, similar to the approach used by Q2D.
@@ C.1 Prompts
Table 6 compares in-context example selection methods. While Q2D-TOP3 uses pseudo-relevant examples, its advantage is limited to R@@10, suggesting that example selection methods do not significantly impact performance. So, we use fixed random examples following Wang et al. (2023).
@@ Appendix F Evaluation Metrics
We evaluate the performance of our retrievers using Recall@@1, Recall@@10, nDCG@@10. Recall@@1 measures the proportion of queries for which the correct passage is ranked first in the retrieved list. Recall@@10 extends this by measuring the proportion of queries for which the correct passage appears in the top 10 retrieved passages. It reflects the model’s ability to identify relevant passages within a broader set of candidates. nDCG@@10 (Normalized Discounted Cumulative Gain at 10) considers the position of relevant passages, giving higher weight to passages ranked closer to the top.
@@ Appendix I Case Studies
Other baseline methods struggle to retrieve the target passage due to vocabulary mismatches between the query and the target passage (Table 9, 12) or because the expanded query includes irrelevant information which may incur hallucination problems mentioned in Introduction (Table 10, 11). However, GuRE generates a query identical to the target passage (Table 13).